# TBRD: TESLA Authenticated UAS Broadcast Remote ID

### v1

@@ III-B System Overview
This work aims to design a system that meets Remote ID requirements while adding protections against the attacks defined by our threat model. Our system comprises three major components: the UAS (with either an integrated RID transponder or an add-on module), a mobile phone with TEE, and a trusted USS verification server. Leveraging the TRBD protocol, these components enable UAS pilots to transmit authenticated Remote ID and observers to verify broadcasts. Our system design is depicted in Figure 2.
@@
Figure 2: TBRD System Diagram. Components, including the UAS, mobile device, TEE, and USS, and how they interact to secure Remote ID broadcasts.
@@
Communication between the system components will occur according to the following procedure. First, a UAS pilot registers as an operator with the FAA and is assigned an Operator ID, which ties any UAS operations to the individual operator. Furthermore, the operator registers their UAS to receive a unique UAS ID, which identifies the individual UAS. Before the start of the first UAS mission (i.e., a new UAS), the UAS pilot registers the UAS ID (UAS serial number or other number assigned by the manufacturer) with the USS (UAS Service Supplier). The UAS pilot’s mobile device is linked to the UAS ID and Operator ID during registration. The commissioning of the new UAS is accomplished over a secure channel. The mobile device becomes, in essence, a physical token that allows the UAS pilot to inform the USS of future UAS missions. Before a UAS mission begins, the pilot configures the UAS’s Operator ID, the UAS ID (which can be the UAS’s serial number or another number assigned by the manufacturer), and the intended flight duration (i.e., the validity period). The mobile device uses the trusted execution environment to generate all the interval keys, $K_{1},...K_{n-1}$, that will be used to sign TBRD broadcast messages, and the key commitment, $K_{0}$, that will be sent to the USS. The key commitment and validity period are transmitted to the USS via a secure communications channel. The USS only accepts mission plans from the mobile devices provisioned during account commission. The interval keys, $K_{1},...K_{n-1}$, are uploaded to the UAS via a secure communication channel. In practice, this could be a secure Bluetooth communication link. The mission start time (defined by the timestamp of the first TBRD message broadcast) is transmitted to the USS. The UAS broadcasts TBRD messages until all interval keys have been exhausted. After all the interval keys are exhausted, the UAS broadcasts in a non-authenticated mode. During the flight, an observer captures at least two TBRD message broadcasts. The interval key disclosure is used to authenticate any previously broadcast message. The observer queries the USS via a secure communications channel to check the validity of the root key, Operator ID, and UAS ID.
III-C TESLA-Based Remote ID Authentication
@@
Figure 3: TESLA key chain generation. Keys are generated in reverse order of key disclosure and message authentication.
@@
The TESLA protocol operates as follows. Figure 3 depicts the process of creating and using keys to generate asymmetric authentication properties, using symmetric keys. The sender divides anticipated transmission time into $n+1$ intervals. The intervals are defined by the index set $I=\{0,1,2,...n\}$. The sender and receiver are loosely synchronized, and the receiver is aware of the time intervals in advance. The sender then generates a set of keys using a one-way keychain function. We define the set of keys $K=\{K_{i}|i\in I\}$. As a result of the one-way key generation function, knowledge of $K_{i-1}$ does not allow the computation or recovery of $K_{i}$. Additionally, the keys are generated in reverse order of key use, where $K_{n}$ is generated first but used last. To prevent the reuse of keys (i.e., using $K_{I}$ multiple times), a pseudo-random function is used to generate $K^{\prime}_{i}$, which will be used to generate a message authentication code (MAC). The sender can now begin to generate broadcast messages. If the sender wishes to send a message $M_{j}$, where $j$ is the current interval index, $i$ is the corresponding key interval, and $d$ is an adjustable disclosure delay, they construct packet $P_{j}$ which is outlined by equation 1.
$P_{j}=\{M_{j}||MAC(K^{\prime}_{i},M_{j}||K_{i-d})\}$ (1)
@@
The disclosure of the previous interval key $K_{i-d}$ with the broadcast at the current interval $j$ allows a receiver to authenticate all previous messages. However, the receiver must wait until the next interval to authenticate the message broadcast in the current interval. A receiver can also generate $K_{0}$, to validate that any interval key, $K_{i}$, is a valid member of the set $K$. In this manner, the value $K_{0}$ becomes a token to validate the commitment to the set of keys from the one-way keychain. If either the validation of the message itself fails or the interval key is not a member of the one-way keychain, the message cannot be authenticated.
@@
For TESLA to operate correctly, several conditions must be met. First, as mentioned previously, the sending and receiving entities must have loose time synchronization. In other words, the receiver must know the upper bound of the difference between the sender’s and receiver’s clocks. The key interval must be chosen such that it takes into account this error. Second, the receiver must know the key disclosure schedule, including the interval duration, $T_{int}$, key disclosure delay $d$, and a key commitment to the keychain, $K_{i}$. Section IV demonstrates how TBRD fulfills these requirements by utilizing tightly synchronized GPS time, Wi-Fi beacon broadcasts, standardized key generation via the TEE of a mobile device, and verification of the key commitment through a trusted third party.
@@ III-E Observer Verification via USS
To satisfy the remaining security requirements, we need to provide a mechanism for an observer to verify the keychain in use. Additionally, we must protect from an attacker simply replaying an old mission. In addition to the attacks outlined in III-A, there are many other scenarios where an attacker may want an observer to receive conflicting or incorrect Remote ID information. Therefore, we must provide additional system functionality that connects the intended mission with the observer to verify flight data. We propose the following.
@@
Once the mobile device has generated the set of keys, the key commitment value, and mission start, they are published to a trusted server. Once the mission is complete, the mobile device sends the mission end time. The ASTM standard proposes the use of a UAV Service Supplier (USS) to facilitate the coordination and distribution of network Remote ID  [16]. In network Remote ID, the operator publishes Remote ID information to a USS via an active internet connection from a ground control station (i.e., cellular or Wi-Fi). We adopt this concept of the trusted USS into our system design. When an observer receives a TBRD-authenticated broadcast from a UAS, the observer first verifies the authenticity of the received packet using the follow-on key disclosure. The USS can then provide the mission’s pre-disclosed key commitment value, start, and end time. We assume that all these transactions occur over cryptographically secure connections. If the observer receives a message but cannot immediately query the USS, this verification process can be accomplished at a later date. Additional assumptions and specific verification algorithms will be further discussed in section IV.
IV Implementation
@@
We implemented our design by leveraging the existing protocols defined by ASTM F3411-22a and the open-source code base OpenDroneID  [17]. OpenDroneID is a C-based library that encodes/decodes OpenDroneID messages, which can be transmitted/received over a wireless interface, such as Wi-Fi. To facilitate rapid prototyping and easier integration with existing simulation infrastructure, we implemented our system using Python. Our source code for a proof-of-concept implementation can be found here. The implementation includes a mobile application, a Linux-based transmitter, receiver, and USS server. A more detailed description of our implementation can be found in the Appendix.
@@
The ASTM standard offers four broadcast transport mechanisms for broadcast Remote ID, primarily driven by the available payload capacity. Those options are Bluetooth Legacy (4.x), Bluetooth 5.x Long Range, Wi-Fi Neighbor Awareness Network (NAN), and Wi-Fi Beacon (vendor-specific information element in the SSID beacon frame). We choose to implement broadcast via Wi-Fi Beacon. As discussed in the previous section V, special attention must be paid to selecting a wireless broadcast transport mechanism when utilizing TESLA authentication. Wi-Fi beacons allowed for a key interval large enough to comply with ASTM F3411-22a and maintain security guarantees of the TESLA protocol. We have initially validated that our system does comply with ASTM F3411-22a utilizing the open source OpenDroneID Android application  [25]. Figure 4 shows our TBRD WiFi broadcast successfully received and parsed by the application, demonstrating compatibility with the standard.
@@
Figure 4: OpenDroneID application displaying the authentication messages transmitted by our proposed system, demonstrating successful message reception and validation on standard Remote ID monitoring devices.
IV-A Mobile Application and TEE
@@
The mobile application serves as the primary interface for pilots to initiate and manage UAS missions, while also facilitating secure key generation by utilizing the TEE on the mobile device. [26]
@@
When the entire keychain is generated, it is exported from the TEE and uploaded to the UAS via a secure upload mechanism. Next, the app creates a file containing the UAS ID, Operator, ID, mission start time, mission end time, and the $n+1$ key. This is sent to the USS securely for TBRD message verification by any observer. In our proof-of-concept implementation, this step is simulated by manually transferring the key file to the TBRD transmitter application. A depiction of the mobile application user interface can be found in the Appendix.
IV-B TBRD Broadcast Transmitter
@@
Before flight, the interval keys are uploaded to the UAS via a secure communications channel. Implementing a secure upload of keys from the mobile application to the UAS is beyond the scope of this work. We assume that there are numerous ways to execute this step securely and efficiently. The interval keys $(K_{i})$ and the root Key $(K_{0})$ are saved in the transmitter memory. The number of keys and key length must be chosen considering the storage space available on the transmitter.
@@
$(i\|\text{HMAC(Authentication Payload},K_{i})\|K_{i-d})$ (4)
@@
Because we opted to send our TBRD messages within the vendor information element of a broadcast Wi-Fi beacon, special care must be taken to ensure this implementation does not violate the security requirements of the TESLA protocol. Specifically, we need to ensure that once the broadcast beacon message has been crafted, it can be sent within the intended interval. However, a wireless network defined by IEEE 802.11 may prevent the beacon’s broadcast within the interval. Wi-Fi networks implement a carrier sense multiple access with collision avoidance (CSMA/CA) feature, which requires the transmitting station to verify that the channel is idle before transmission [27]. If the channel is busy, the transmitting station must wait a specified amount of time before attempting transmission. Therefore, the delay may cause the transmission to occur outside the intended interval. The transmission of a message outside its intended interval has profound security implications. For example, a system implements a key disclosure delay of $d=1$. Message $M_{4}$ is a message meant for transmission in interval $i=4$. If $M_{4}$ is delayed and sent in interval $i=5$ along with Message $M_{5}$, the disclosure of the key used to authenticate $M_{4}$ occurs in the same interval of key disclosure. This would violate the security guarantees offered by the delayed key disclosure mechanisms. As a result, we must be ready to discard and prevent the transmission of any packets outside the intended interval period.
Controlling the transmission of Wi-Fi beacons within a software implementation of the TBRD protocol requires implementing additional safeguards to manage a potentially dynamic wireless channel effectively. One approach would be to build custom Wi-Fi firmware to prevent beacons from being broadcast outside their intended interval. However, we chose an approach that defined a permitted broadcast period within the key interval. This approach also considers the TESLA protocol enhancements from [28], which address the maximum clock skew expected from GPS synchronization errors.
Figure 5: Definition of a permissible transmission window within each key interval. This is the period where Wi-Fi beacon broadcasts are authorized to ensure a balance between system functionality and security requirements.
@@ IV-C TBRD Receiver and Verification
Figure 6: Key verification workflow showing how the observer authenticates received messages through subsequent key disclosure, USS coordination for key-commitment validation, and temporal checks to confirm that keys remained secret during their usage intervals. The $K$ in the orange box represents the symmetric key used to encrypt the message package $P$, while the $K$ in the green box denotes the key disclosed in interval $i$ to enable verification of the corresponding earlier message.
Figure 6 uses a specific scenario where the observer first observes a message transmitted at interval 4. The interval period $T$ equals one second, and the disclosure delay is 1.
  1. 1.
    The observer receives message $M_{4}$ at some time $t$ and waits for the receipt of $M_{5}$. Once received, it uses the key disclosed in $M_{5}$ to validate the authenticity of the HMAC broadcast in $M_{4}$.
  2. 2.
@@
IV-D USS Server
The verification server, or USS, stores mission-specific information to validate TBRD messages. The exact implementation of the server could vary significantly in a production environment. However, the basic functional requirements are that it must be able to communicate with authorized users, store the mission data, and respond to mission data queries from receivers. Our implementation utilizes a simple Python dictionary to store mission data and communicates with the transmitter and receiver using ZeroMQ and Google Protocol Buffers  [29, 30].
@@ IV-D USS Server
IV-E Swarm Simulation Testbed
@@
To evaluate the robustness of the TBRD protocol and our implementation, we simulated an autonomous UAS swarm utilizing. In the swarm simulation, each UAS relies on the location information broadcast by Remote ID to make collision avoidance decisions while executing a pre-defined mission. We demonstrate several attacks on broadcast transmission and evaluate the extent to which TBRD can protect from these attacks.
@@
We leveraged an internal UAS swarm simulation testbed (name anonymized for review) to enable rapid simulation and prototyping of autonomous UAS swarm algorithms. The software architecture is portrayed in Figure 7. At the heart of the system is the Gazebo robotics simulator. The Gazebo platform can simulate physical environments, real-world sensors, and communication transport mechanisms. The testbed utilizes Docker to simulate a swarm network, connecting a set of containerized ArduPilot flight controllers [31]. ArduPilot is an open-source autopilot solution that can simulate various types of autonomous vehicles  [32]. A current limitation of the testbed platform is the inability to simulate a wireless communications channel. As a result, we chose to implement communications over the Docker network where the contents of the WiFi beacons were broadcast via UDP. While this does not perfectly emulate a wireless channel that must comply with the 802.11 protocol, the connectionless nature of UDP broadcast allows us to evaluate the critical features of the TBRD system. With these combined features, the testbed can emulate a fully decentralized UAS swarm architecture capable of executing complex mission scenarios.
Figure 7: UAS Swarm Simulation Testbed System Architecture. The Gazebo robotics simulator is used in conjunction with a variable number of Docker containers to facilitate high-fidelity simulation of a decentralized UAS swarm.
Collision Avoidance Algorithm: One challenge that autonomous UAS swarms must overcome is executing a given mission or task without collision between individual agents of the swarm. The ability to sense the surrounding environment, make decisions, and autonomously execute a mission is a key requirement to enable decentralized control of a UAS swarm. In the perception-decision-action loop that drives a large-scale UAS swarm, an inability to perceive surrounding UAS efficiently will significantly reduce the swarm’s effectiveness  [33]. There are many technological approaches to the initial steps of perception. These vary from visual detection, by either direct or passive means, audio, and infrared sensors  [34]. Furthermore, several works have implemented solutions that center around communication-based ranging to provide adequate situational awareness  [35, 36, 37, 38, 39]. These solutions rely on some form of active or passive communications channel to determine distance or relative location between individual UASs. While the abovementioned solutions afford various strengths and weaknesses, none protect integrity. In contrast, TBRD provides a robust means of broadcasting position, with an added layer of authentication to ensure only trusted locations are used for collision avoidance calculations.
@@
Once the locations of surrounding agents in the swarm have been collected, some form of algorithm or procedure must process that information to decide how to act. Evaluating the relative performance of differing collision avoidance algorithms is beyond the scope of this investigation. Instead, we aimed to evaluate the relative safeguards against attacks against communications-based ranging when using TBRD. As such, we evaluated two potential collision avoidance algorithms to provide decentralized decision-making capabilities to our swarm simulation: optimal reciprocal collision avoidance (ORCA)  [40], and Global to Local Autonomy Synthesis (GLAS)  [41]. Based on the availability of code implementation, we used ORCA as our collision avoidance decision algorithm  [42]. ORCA uses the concept of velocity obstacles to derive collision-free conditions for a set of system parameters [40]. While ORCA initially provided collision avoidance capabilities in two dimensions, it has continued to evolve to execute collision avoidance in three-dimensional space  [43]. However, we have restricted our evaluation scenario to exploring collision avoidance in 2-dimensions. Thus, the altitude of each UAS in the swarm is limited to a static height, allowing us to conduct our evaluation. The collision avoidance algorithm calculates velocities to avoid collision at a configurable time step value, $t$. At each time step $t$, a collision avoidance vector is calculated such that collision is guaranteed for the next $\tau$ seconds. Furthermore, a radius $r$ is chosen, which can be viewed as a safety margin that is accounted for when computing collision avoidance vectors.
@@
Experimental Setup: Our experiment consisted of a simulation of 4 UASs, each beginning their mission on the corner of a 40-meter by 40-meter square. Each UAS is instructed to take off, fly to an altitude of 5 meters, and then fly to the opposing corner of the square, landing upon arrival. This high-level mission layer directive is augmented with the ORCA collision avoidance mechanism, whereby each UAS should take into account the location of every other UAS in the swarm and avoid collision. Our implementations of the ORCA algorithm used a timestep of 1 sec, a radius value of 2 meters, and a tautau of 5 seconds. We first developed a baseline mission using Remote ID to share location information. We then simulated two additional scenarios, one using Remote ID with no authentication, and another using TBRD to share location. During the final two scenarios, we also simulated a spoofing attack where a line of UASs was placed along the y-axis. This simulates a scenario where an attacker has knowledge of the intended mission scenario and is within close proximity to inject fake Remote ID broadcasts.
Figure 8: TBRD Evaluation Scenarios Results. This figure illustrates how the UAS swarm moves without Remote ID, with no attack, and with Remote ID authentication spoofing a line of UASs, as well as TBRD authentication with Remote ID spoofing UASs.
@@
V Security and Performance Analysis
TABLE I: Comparative Security and Performance Analysis of Remote ID Authentication Approaches

### v2

@@ 3.2. System Overview
Our system is comprised of three major components: the TBRD Transponder (either integrated within the UAS or as an add-on module), a mobile application runing on a device with a TEE, and a trusted USS verification server. Leveraging the TRBD protocol, these components enable UAS pilots to transmit authenticated Remote ID and observers the ability to verify broadcasts. Our system design is depicted in Figure 2.
@@
Figure 2. TBRD System Diagram. Components, including the UAS, mobile device, TEE, and USS, and how they interact to secure Remote ID broadcasts.: TBRD System Diagram
@@
Communication between the system components will occur according to the following procedure. First, a UAS pilot registers as an operator with the FAA and is assigned an Operator ID, which ties any UAS operations to the individual operator. Furthermore, the operator registers their UAS to receive a unique UAS ID, which identifies the individual UAS. Before the start of the first UAS mission (i.e., a new UAS), the UAS pilot registers the UAS ID (UAS serial number or other number assigned by the manufacturer) with the USS. The UAS pilot’s mobile device is linked to the UAS ID and Operator ID during registration. The commissioning of the new UAS is accomplished over a secure channel. The mobile device becomes, in essence, a physical token that allows the UAS pilot to inform the USS of future UAS missions. Before a UAS mission begins, the pilot configures the UAS’s Operator ID, the UAS ID (which can be the UAS’s serial number or another number assigned by the manufacturer), and the intended flight duration (i.e., the validity period). The mobile device uses the trusted execution environment to generate all the interval keys, $K_{1},...K_{n-1}$, that will be used to sign TBRD broadcast messages, and the key commitment, $K_{0}$, that will be sent to the USS. The key commitment and validity period are transmitted to the USS via a secure communications channel. The USS only accepts mission plans from the mobile devices provisioned during account commission. The interval keys, $K_{1},...K_{n-1}$, are uploaded to the UAS via a secure communication channel. In practice, this could be a secure Bluetooth communication link. The mission start time (defined by the timestamp of the first TBRD message broadcast) is transmitted to the USS. The UAS broadcasts TBRD messages until all interval keys have been exhausted. After all the interval keys are exhausted, the UAS broadcasts in a non-authenticated mode. During the flight, an observer captures at least two TBRD message broadcasts. The interval key disclosure is used to authenticate any previously broadcast message. The observer queries the USS via a secure communications channel to check the validity of the root key, Operator ID, and UAS ID.
3.3. TESLA-Based Remote ID Authentication
@@
Figure 3. TESLA key chain generation. Keys are generated in reverse order of key disclosure and message authentication.: TESLA Background
@@
The TESLA protocol achieves asymmetric authentication properties, using symmetric authentication keys. Figure 3 depicts the process of key generation. The sender divides anticipated transmission time into $n+1$ intervals. The intervals are defined by the index set $I=\{0,1,2,...n\}$. The sender and receiver are loosely time synchronized, and the receiver is aware of the time intervals in advance. The sender then generates a set of keys using a one-way keychain function. We define the set of keys $K=\{K_{i}|i\in I\}$. As a result of the one-way key generation function, knowledge of $K_{i-1}$ does not allow the computation or recovery of $K_{i}$. Additionally, the keys are generated in reverse order of key use, where $K_{n}$ is generated first but used last. To prevent the reuse of keys (i.e., using $K_{I}$ multiple times), a pseudo-random function is used to generate $K^{\prime}_{i}$, which will be used to generate a message authentication code (MAC). The sender can now begin to generate broadcast messages. The sender generates a message $M_{j}$, where $j$ is the current interval index, $i$ is the corresponding key interval, and $d$ is an adjustable disclosure delay. The constructed packet $P_{j}$ is defined by equation 1.
(1) $P_{j}=\{M_{j}||MAC(K^{\prime}_{i},M_{j}||K_{i-d})\}$
@@
The disclosure of the previous interval key $K_{i-d}$ along with the message $M_{j}$, allows a receiver to authenticate all previous messages. However, the receiver must wait until the next interval to authenticate the message broadcast in the current interval. A receiver can also generate $K_{0}$, to validate that any interval key, $K_{i}$, is a valid member of the set $K$. In this manner, the value $K_{0}$ becomes a token to validate the commitment to the set of keys from the one-way keychain. If either the validation of the message itself fails or the interval key is not a member of the one-way keychain, the message cannot be authenticated.
@@
For TESLA to operate correctly, several conditions must be met. First, as mentioned previously, the sending and receiving entities must have loose time synchronization. In other words, the receiver must know the upper bound of the difference between the sender’s and receiver’s clocks. The key interval must be chosen such that it takes into account this error. Second, the receiver must know the key disclosure schedule, including the interval duration, $T_{int}$, key disclosure delay $d$, and a key commitment to the keychain, $K_{i}$. Section 4 demonstrates how TBRD fulfills these requirements by utilizing tightly synchronized GPS time, Wi-Fi beacon broadcasts, standardized key generation via the TEE of a mobile device, and verification of the key commitment through a trusted third party.
@@ 3.5. Observer Verification via USS
To satisfy the remaining security requirements, we need to provide a mechanism for an observer to verify the keychain in use. Additionally, we must protect our system from an attacker who simply replays broadcasts from an old mission. In addition to the attacks outlined in 3.1, there are many other scenarios where an attacker may want an observer to receive conflicting or incorrect Remote ID information. Therefore, we must provide additional system functionality that connects the intended mission with the observer to verify flight data. We propose the following.
@@
Once the mobile device has generated the set of keys, the key commitment value, and mission start, they are published to a trusted server. Once the mission is complete, the mobile device sends the mission end time. The ASTM standard proposes the use of a UAV Service Supplier (USS) to facilitate the coordination and distribution of network Remote ID  (f38_committee_specification_2022). In network Remote ID, the operator publishes Remote ID information to a USS via a persistent internet connection from a ground control station (i.e., cellular or Wi-Fi). We adopt this concept of the trusted USS into our system design, however the internet connection no longer must be persistent. When an observer receives a TBRD-authenticated broadcast from a UAS, the observer first verifies the authenticity of the received packet using the follow-on key disclosure. The USS can then provide the mission’s pre-disclosed key commitment value, start, and end time. We assume that all these transactions occur over cryptographically secure connections. If the observer receives a message but cannot immediately query the USS, this verification process can be accomplished at a later date. Additional assumptions and specific verification algorithms will be further discussed in section 4.
4. Implementation
@@
We implemented our design by leveraging the existing protocols defined by ASTM F3411-22a and the open-source code base OpenDroneID  (open_drone_id_open_2025). OpenDroneID is a C-based library that encodes and decodes OpenDroneID messages, which can be transmitted/received over a wireless interface, such as Wi-Fi. To facilitate rapid prototyping and easier integration with existing simulation infrastructure, we implemented our system using Python. Our source code for a proof-of-concept implementation can be found here. The implementation includes a mobile application, a Linux-based transmitter, receiver, and USS server.
@@
The ASTM standard offers four broadcast transport mechanisms for broadcast Remote ID, primarily driven by the available payload capacity. Those options are Bluetooth Legacy (4.x), Bluetooth 5.x Long Range, Wi-Fi Neighbor Awareness Network (NAN), and Wi-Fi Beacon (vendor-specific information element in the SSID beacon frame). We choose to implement broadcast via Wi-Fi Beacon. Special attention must be paid to selecting a wireless broadcast transport mechanism when utilizing TESLA authentication. Wi-Fi beacons allowed for a key interval large enough to comply with ASTM F3411-22a and maintain security guarantees of the TESLA protocol. We have validated that our system does comply with ASTM F3411-22a utilizing the open source OpenDroneID Android application  (open_drone_id_opendroneidreceiver-android_2025). Figure 4 shows our TBRD WiFi broadcast successfully received and parsed by the application, demonstrating compatibility with the standard.
@@
Figure 4. OpenDroneID application displaying the authentication messages transmitted by our proposed system, demonstrating successful message reception and validation on standard Remote ID monitoring devices.: DroneScanApp verification
4.1. Mobile Application and TEE
@@
The mobile application serves as the primary interface for pilots to initiate and manage UAS missions, while also facilitating secure key generation by utilizing the TEE on the mobile device. (android_android_2025)
@@
When the entire keychain is generated, it is exported from the TEE and uploaded to the UAS via a secure upload mechanism. Next, the app creates a file containing the UAS ID, Operator, ID, mission start time, mission end time, and the $n+1$ key. This is sent to the USS securely for TBRD message verification by any observer. In our proof-of-concept implementation, this step is simulated by manually transferring the key file to the TBRD transmitter application.
4.2. TBRD Broadcast Transmitter
@@
Before flight, the interval keys are uploaded to the UAS via a secure communications channel. Implementing a secure upload of keys from the mobile application to the UAS is beyond the scope of this work. There are numerous ways to execute this step securely and efficiently. The interval keys $(K_{i})$ and the root Key $(K_{0})$ are saved in the transmitter memory. The number of keys and key length must be chosen considering the storage space available on the transmitter.
@@
(4) $(i\|\text{HMAC(Authentication Payload},K_{i})\|K_{i-d})$
@@
Because we opted to send our TBRD messages within the vendor information element of a broadcast Wi-Fi beacon, special care must be taken to ensure this implementation does not violate the security requirements of the TESLA protocol. Specifically, we need to ensure that once the broadcast beacon message has been crafted, it can be sent within the intended interval. However, a wireless network defined by IEEE 802.11 may prevent the beacon’s broadcast within the interval. Wi-Fi networks implement a carrier sense multiple access with collision avoidance (CSMA/CA) feature, which requires the transmitting station to verify that the channel is idle before transmission (ieee_ieee_2025). If the channel is busy, the transmitting station must wait a specified amount of time before attempting transmission. Therefore, the delay may cause the transmission to occur outside the intended interval. The transmission of a message outside its intended interval has profound security implications. For example, a system implements a key disclosure delay of $d=1$. Message $M_{4}$ is a message meant for transmission in interval $i=4$. If $M_{4}$ is delayed and sent in interval $i=5$ along with Message $M_{5}$, the disclosure of the key used to authenticate $M_{4}$ occurs in the same interval of key disclosure. This would violate the security guarantees offered by the delayed key disclosure mechanisms. As a result, we must be ready to discard and prevent the transmission of any packets outside the intended interval period.
Controlling the transmission of Wi-Fi beacons within a software implementation of the TBRD protocol requires implementing additional safeguards to manage a potentially dynamic wireless channel effectively. One approach would be to build custom Wi-Fi firmware to prevent beacons from being broadcast outside their intended interval. However, we chose an approach that defined a permitted broadcast period within the key interval. This approach also considers the TESLA protocol enhancements from (jahanian_analysis_2015), which address the maximum clock skew expected from GPS synchronization errors.
Figure 5. Definition of a permissible transmission window within each key interval. This is the period where Wi-Fi beacon broadcasts are authorized to ensure a balance between system functionality and security requirements.: Permitted Tx Time
@@ 4.3. TBRD Receiver
Figure 6. Key verification workflow showing how the observer authenticates received messages through subsequent key disclosure, USS coordination for key-commitment validation, and temporal checks to confirm that keys remained secret during their usage intervals.: Verification process
Figure 6 demonstrates a specific scenario where the observer receives a message transmitted at intervals 1 through 4. The interval period $T$ equals one second, and the disclosure delay is 1.
  1. (1)
    The observer receives message $M_{3}$ at some time $t$ and waits for the receipt of $M_{4}$. Once received, it uses the key disclosed in $M_{4}$ to validate the authenticity of the HMAC broadcast in $M_{3}$.
  2. (2)
    USS Verification:
    1. (a)
@@
4.4. USS Server
The verification server, or USS, stores mission-specific information to validate TBRD messages. The exact implementation of the server could vary significantly in a production environment. However, the basic functional requirements are that it must be able to communicate with authorized users, store the mission data, and respond to mission data queries from receivers. Our implementation utilizes a simple Python dictionary to store mission data and communicates with the transmitter and receiver using ZeroMQ and Google Protocol Buffers  (google_protocol_2025; zeromq_zeromq_2025).
@@ 4.4. USS Server
5. Security and Performance Analysis
5.1. Swarm Simulation Testbed
@@
To evaluate the robustness of the TBRD protocol and our implementation, we simulated an autonomous UAS swarm. In the swarm simulation, each UAS relies on the location information broadcast by Remote ID to make collision avoidance decisions while executing a pre-defined mission. We demonstrate several attacks on broadcast transmission and evaluate the extent to which TBRD can protect from these attacks.
@@
We leveraged an internal UAS swarm simulation testbed (name anonymized for review) to enable rapid simulation and prototyping of autonomous UAS swarm algorithms. At the heart of the system is the Gazebo robotics simulator. The Gazebo platform can simulate physical environments, real-world sensors, and communication transport mechanisms. The testbed utilizes Docker to simulate a swarm network, connecting a set of containerized ArduPilot flight controllers (docker_docker_2025). ArduPilot is an open-source autopilot solution that can simulate various types of autonomous vehicles  (ardupilot_ardupilot_2025). A current limitation of the testbed platform is the inability to simulate a wireless communications channel. As a result, we chose to implement communications over the Docker network where the contents of the WiFi beacons were broadcast via UDP. While this does not perfectly emulate a wireless channel that must comply with the 802.11 protocol, the connectionless nature of UDP broadcast allows us to evaluate the critical features of the TBRD system. With these combined features, the testbed can emulate a fully decentralized UAS swarm architecture capable of executing complex mission scenarios.
Figure 7. This figure demonstrates the positive impact of TBRD Authentication on collision avoidance for swarms. The first graph indicates the path traveled by four drones during a baseline scenario of using unauthenticated Remote ID. The second graph illustrates how the same four drones will adjust their paths to avoid spoofed drones. The last graph shows that the four drones can successfully detect spoofed drones when using TBRD Authentication and complete their mission without interruption. : TBRD Evaluation
Collision Avoidance Algorithm: One challenge that autonomous UAS swarms must overcome is executing a given mission or task without collision between individual agents of the swarm. TBRD provides a robust means of broadcasting position, with an added layer of authentication to ensure only trusted locations are used for collision avoidance calculations.
@@
Once the locations of surrounding agents in the swarm have been collected, some form of algorithm or procedure must process that information to decide how to act. Evaluating the relative performance of differing collision avoidance algorithms is beyond the scope of this investigation. Instead, we aimed to evaluate the relative safeguards against attacks against communications-based ranging when using TBRD. As such, we evaluated one such potential collision avoidance algorithm that provides decentralized decision-making capabilities to our swarm simulation: optimal reciprocal collision avoidance (ORCA)  (siciliano_reciprocal_2011). The availability of code implementation weighed heavily on our decision to use ORCA as our collision avoidance decision algorithm  (nazecic-andrlon_muonpyorca_2025). ORCA uses the concept of velocity obstacles to derive collision-free conditions for a set of system parameters (siciliano_reciprocal_2011).
@@
Experimental Setup: Our experiment consisted of a simulation of 4 UASs, each beginning their mission on the corner of a 40-meter by 40-meter square. Each UAS is instructed to take off, fly to an altitude of 5 meters, and then fly to the opposing corner of the square, landing upon arrival. This high-level mission layer directive is augmented with the ORCA collision avoidance mechanism, whereby each UAS should take into account the location of every other UAS in the swarm and avoid collision throughout the mission. Our implementations of the ORCA algorithm used a timestep of 1 sec, a radius value of 2 meters, and a $\tau$ of 5 seconds. We first developed a baseline mission using Remote ID to share location information. We then simulated two additional scenarios, one using Remote ID with no authentication, and another using TBRD to share location. During the final two scenarios, we also simulated a spoofing attack where a line of UASs was placed along the y-axis. This simulates a scenario where an attacker has knowledge of the intended mission scenario and is within close proximity to inject fake Remote ID broadcasts.
@@
5.2. Spoofing-Replay-Timing Attack Resilience
Table 1. Comparative Security and Performance Analysis of Remote ID Authentication Approaches